a certain mass of a gas is expanded from (1l,10 atm) to (4l,5atm) against a constant external pressure of 1 atm. If initial temperature of a gas is300 k and the heat capacity of the process is 50 j/c. The enthalpy is ?

Grade:11

3 Answers

Susmita
5 years ago
Enthalpy change isCpdTdT is change in temperature.Given P1=10atm,V1=1l,T1=300k,P2=5atm,V2=4l.P1V1=RT1 and P2V2=RT2So,T
Susmita
5 years ago
T2=P2V2T1/P1V1=600KSo
Susmita
5 years ago
Enthalpy change =CpdT=50×(600-300)=15000J=15kJ
..................................................................................................................................